A man is known to speak truth 3 out of 4 times.He throws adie and reports it is a six. Find the probability that it is actually a six.

Grade:12

2 Answers

hari krishn
10 years ago

(3/4)*(1/6) = (1/8)

hari krishn
10 years ago

probability of speaking truth is 0.75


and that of speaking false is 0.25


probability of coming 6=1/6

probabiliy that it is 6 and he say the truth = .75*1/6

probability that it is not 6 and he say it as 6 = .25*5/6

Probability that it is actually 6=(0.75*1/6) / {(0.75*1/6) + (0.25*5/6)}


                                             =3/8